# UniCode: A Framework for Generating High-Quality Competitive Coding Problems

### v1

@@ B.0.1 Setup and Notation
B.0.2 Systematic Bias (Identification and De-biasing)
@@
B.0.3 Random Error (Sampling Variability)
@@
By concavity of x(1x)x(1-x), equation 4 $\leq$ equation 6, so using equation 6 is conservative when the split is fixed.
@@
where $\underline{\mu},\overline{\mu}]$ is a $95\%$ CI for $\mu$. If only bounds are known ($\alpha\in[0,\alpha_{\max}]$, $q_{e}\in[q_{\min},q_{\max}]$), combine equation 3 with $\underline{\mu},\overline{\mu}]$ to obtain a conservative CI for $p$:
@@
B.0.4 Total Error Bound
@@
When $n$ is large, the $O(n^{-1/2})$ term vanishes and the total error is controlled by the bias ceiling $\alpha$. If $\alpha$ (and $q_{e}$) are known, report the de-biased estimate equation 2 with CI equation 8; this both removes the bias and shrinks the CI.
B.0.5 Stratified (Tag-wise) Contamination
@@ B.0.6 Numerical Illustration
B.0.7 Practical Safeguards
  • Audit and bound $\alpha$. Spot-check a random subsample to obtain an empirical upper bound $\alpha_{\max}$ with binomial CIs; report $p$ using equation 9.
@@ B.0.7 Practical Safeguards
  • De-bias when possible. If $(\alpha,q_{e})$ are fixed by design (e.g., known faulty items), publish the de-biased estimate equation 2 and its CI equation 8.
@@
  • Robust reporting. Alongside $\hat{\mu}$, report (i) de-biased $\tilde{p}$, (ii) contamination-aware CIs, and (iii) sensitivity bands under $(\alpha,q_{e})$ ranges as in equation 9.
@@
Appendix C Algorithm Capacity Analysis
We analyze how algorithms of different time complexities behave under constrained time limits. Consider a programming competition problem requiring:
@@ C.0.1 Capacity Analysis
Appendix D Code-tag Distribution
Figure 8: Distribution of tags, sub-tags, and skills in UniCode dataset.
Appendix E Human Rating Website
To ensure the quality and novelty of the problems we generate, we developed a website for human evaluation. The website features a clear and intuitive layout. On the left side of the screen, the seed question and the newly generated question are displayed alongside their test cases, allowing human raters to compare and evaluate them directly. Below the questions, rating criteria are provided, including "Solvability," "Novelty," and "New Problem Variant Type." This feedback mechanism is essential for refining our framework and enhancing the quality of the generated problems.
Figure 9: Human rating website.
Appendix F Example Problems
F.1 Seed Problem 1
Knowing the devilish nature of Suraj, Ashi is concerned about the Chef and has asked all of Chef’s town people to help. The distance from Chef’s town to Byteland through the the path of tasty dishes is $X$ units. They have the location where the magic dishes are and how many people are required to eat it completely. Anyone who eats a dish would go to a long sleep and won’t be able to continue. They have the information about the tribal clans that live along the the path of tasty dishes who can be of real help in this journey.
The journey Chef and his friends can be described as follows: There is a total of $B$ dishes on the path of tasty dishes. Each dish is located at some distance from Chef’s town denoted by $x_{i}$ for the $i^{th}$ dish ($x_{i-1}<x_{i}$). To minimize the number of friends Chef has to leave behind, all of them have decided that exactly $y_{i}$ of them will eat the $i^{th}$ dish, which is the required number of people needed to finish it completely. Also, there are a total of $C$ tribal chef clans, each with their own population and location on the path that Chef and his friends will meet on their way to Byteland. They know that for some clan (say $i$), they are located at a distance of $p_{i}$ ($p_{i-1}<p_{i}$) from Chef’s town with a population of $r_{i}$. And if a group of at least $q_{i}$ men approaches them, they would be able to convince them to join their forces against Suraj.
Given the information about all this, help the Chef to find out the minimum size of the group (including him and his friends) he should start with to reach Byteland and get back Ashi’s gift from Suraj.
[…]
The first line of the input contains an integer T denoting the number of test cases.Each test case contains three lines which are as follows:First line of each test case contains X, the distance of Byteland from Chef’s town.Next line contains an integer B, the number of dishes on the path of tasty dishes.Then follows B pairs of space separated integers of the form x_i y_i, where x_i y_i are as defined above for the i-th dish.Next line contains an integer C, followed C space separated triplets of integers p_i q_i r_i as defined above.
[…]
For each test case, print the minimum size of the group (including Chef) that is needed to reach Byteland.
Constraints:
  • $1\leq T\leq 10$
  • $1\leq X\leq 10^{9}$
  • $1\leq B\leq 10000$
  • Subproblem 1 (25 points): $C=0$
  • Subproblem 2 (75 points): $1\leq C\leq 10000$
  • $1\leq x_{i}<X,x_{i}<x_{i+1}$
  • $1\leq p_{i}<X,p_{i}<p_{i+1}$
  • $1\leq y_{i}\leq 10^{14}$
  • $1\leq q_{i}\leq 10^{14}$
  • $1\leq r_{i}\leq 10^{14}$
  • All the positions, of the tasty dishes and tribal clans are distinct.
F.2 Seed Problem 2
Problem Statement: Read problems statements in Mandarin Chinese and Russian.
The Baratheons have been ruling in the Seven Kingdoms for many years. They have seen a lot: prosperity and hunger, tranquility and rebellions, live and death. But nevertheless, they still hold the throne. King Joffrey Baratheon’s reign is running now. As a wise man, he honors the history of his family. So, he commanded to build up two monuments, that will remind about some historical periods of the Baratheons. Formally, the Baratheons have been ruling for $N$ years. Every year is described by an integer $A_{i}$, the level of prosperity in $i$-th year. If $i$-th year was a great year, then $A_{i}$ might be a positive integer. Otherwise, if $i$-th year was a horrible year, then $A_{i}$ might be a negative integer. Each historical period can be described as two integers $S$ and $F$, the start and the finish of the period respectively. Of course, $S$ is not greater than $F$ for each period, that we consider in this task. You are to pick two historical periods, but there are some rules:
  • Two periods shouldn’t have common years. I.e. a period [1, 5] has no common years with a period [6, 7];
  • The first period should start earlier than the second one. I.e. a period [1, 5] starts earlier than [6, 7];
  • Two periods shouldn’t be too close to each other. There must be at least $K$ years between the finish of the first period and the start of the second period. I.e. periods [1, 5] and [10, 10] can be chosen in case $K$ equals to 4, while they can’t in case $K$ equals to 5.
The sum of the levels of prosperity in chosen years should be as big as possible. Afterwards, you should report the sum of the levels of prosperity in chosen years to your King. Make everything right, otherwise King Joffrey won’t be merciful!
[…]
The first line of the input contains an integer T denoting the number of test cases.The description of T test cases follows.The first line of each test case contains two integers N and K denoting the length of the Baratheons’ reign and the minimal amount of years between two chosen periods.The second line contains N space-separated integers A_1, A_2, ..., A_N denoting the levels of prosperity in corresponding years.
[…]
For each test case, output a single line containing the required integer.
Constraints:
  • $1\leq T\leq 5$
  • $2\leq N\leq 10^{5}$
  • $0\leq K\leq 10^{5}$
  • $10^{9}\leq A_{i}\leq 10^{9}$ for every $1\leq i\leq N$
  • $K+2\leq N$
F.3 New Problem
Problem Statement: Chef must journey in two phases. Phase 1: Recruitment - There are $N$ districts in Chef’s town, labeled 1…N. District $i$ has $A_{i}$ potential volunteers ($A_{i}$ may be negative: a negative value means the district actually shuns the effort). - Chef may conduct exactly two recruitment campaigns, each on a contiguous interval of districts $[L,R]$. These two intervals must not overlap, and there must be at least $K$ districts between the end of the first and the start of the second. Chef gathers the sum of $A_{i}$ in each chosen interval. His total recruits $H$ is the sum over both intervals (if that sum is negative, he would of course choose intervals giving non-negative sum). Phase 2: Expedition - The path from Chef’s town to Byteland has $B$ magical “dishes” at strictly increasing distances $x_{i}$. To cross dish $i$, exactly $y_{i}$ members of Chef’s party must stop (and thus be lost to sleep). - There are $C$ tribal clans at strictly increasing distances $p_{j}$. Clan $j$ will join Chef’s party and contribute $r_{j}$ people, but only if at the moment Chef arrives at $p_{j}$ his current party size is at least $q_{j}$.
Compute the minimal $G_{0}$ such that Chef can complete Phase 2 alive.
[…]
TFor each test case: N K A_1 A_2 ... A_N B x_1 y_1 ... x_B y_B C p_1 q_1 r_1 ... p_C q_C r_C
[…]
For each test case, print one integer: the minimum G_0.
Constraints:
  • $1\leq T\leq 5$
  • $1\leq N,B,C\leq 10^{5}$
  • $0\leq K<N$
  • $10^{9}\leq A_{i}\leq 10^{9}$
  • $1\leq x_{i}<x_{i+1}\leq 10^{9}$
  • $1\leq y_{i}\leq 10^{14}$
  • $1\leq p_{j}<p_{j+1}\leq 10^{9}$
  • $1\leq q_{j},r_{j}\leq 10^{14}$
F.4 The LLM-Unsolvable Problem
Problem Statement: You have $2n$ heroes, the $i$-th hero has initial power $p_{i}$. There are $m$ special pairs, each pair is two heroes; a special pair must be assigned to the same team. You must split all $2n$ heroes into two teams of size $n$: your team and the other team. After the split, you can perform any number of operations on each hero on your team: in one operation you choose a hero and add $x$ to its power or subtract $x$ from its power (but hero’s power cannot go below 0). Finally, define the MEX of your team’s powers as the smallest non-negative integer not present in the multiset of final powers. Compute the maximum MEX you can achieve by choosing an optimal split.
Input Format:
  • Line 1: three integers $n,m,x$ ($1\leq n\leq 2\cdot 10^{5},0\leq m\leq n,1\leq x\leq 10^{5}$)
  • Line 2: $2n$ integers $p_{1},p_{2},\dots,p_{2n}$ ($0\leq p_{i}\leq 10^{9}$)
  • Next $m$ lines: two integers $a,b$ ($1\leq a<b\leq 2n$), denoting a special pair (heroes $a$ and $b$); pairs are disjoint.
Output Format: One integer—the maximum MEX you can obtain on your team under the best assignment and any number of $\pm x$ operations.
Constraints:
  • Special pairs are disjoint; each hero belongs to at most one pair.
  • You must put exactly $n$ heroes on “your team.”
  • If heroes $a$ and $b$ form a special pair, both must go either to your team or to the other team.
Appendix G Integrated Prompts for Test Case Generation
1I will provide you with a programming problem description, and your task is to generate standardized test input samples using the CYaRon library.
2
3You need to complete the following steps:
@@ Appendix G Integrated Prompts for Test Case Generation
7
8Output format (strictly follow)
9Part 1: Parse Input Constraints
10Specify the input constraints as described in the problem.
11
12Part 2: Code for Test Input Generation
13import cyaron as cy #cyaron version: 0.7.0
14
15def generate_test_input():
16 # set parameters constraints that meet requirements (e.g. 1 <= N <= 300)
17 ...
18 # Generate input using CYaRon
19 input_data = [
20 ...
21 ]
@@
23
24Part 3: Code to Validate Test Input
25def validate_test_input(input_string):
26 # Validation logic
27 return <boolean>
28
29Note:
30- cy.Integer() is not supported; it should be cy.randint.
31- use cy.String.random instead of cy.String
32- The function generate_test_input() should not accept any parameters. You need to generate the input entirely within the function.
33- Generate code following the above format, without starting with ‘‘‘python or similar markers.
Listing 1: Random Input Generator (CYaRon)
1I will provide you with a programming problem description, and your task is to generate adversarial test input samples using the CYaRon library.
2
3You need to complete the following steps:
@@
7
8Output format (strictly follow):
9Part 1: Parse Input Constraints
10Specify the input constraints as described in the problem.
11
12Part 2: Code for Test Input Generation
13import cyaron as cy # cyaron version: 0.7.0
14import random
15
16def generate_test_input():
17 # set parameters constraints that meet requirements (e.g. 1 <= N <= 300)
18 ...
19 # Randomly choose one adversarial strategy
20 strategy = random.choice(["equal_weights", "alternating_large_small", "large_ends"])
21
22 if strategy == "equal_weights":
23 # example: all weights are maximal
24 N = cy.randint(100000, 100000) # fix to worst-case size
25 max_weight = cy.randint(109, 109)
26 k = random.choice([1, N//2, N-1, N])
27 weights = [max_weight] * N
28 elif strategy == "alternating_large_small":
29 N = cy.randint(100000, 100000)
30 max_weight = cy.randint(109, 109)
31 k = random.choice([1, N//2, N-1, N])
32 weights = [max_weight if i%2 else 1 for i in range(N)]
33 else: # large_ends
34 N = cy.randint(100000, 100000)
35 max_weight = cy.randint(109, 109)
36 k = random.choice([1, N//2, N-1, N])
37 weights = [max_weight] + [1]*(N-2) + [max_weight]
38
39 # Build input string
40 input_lines = [f"{N} {k}"] + [str(w) for w in weights]
@@
42
43Part 3: Code to Validate Test Input
44def validate_test_input(input_string):
45 try:
@@
47 N_k = lines[0].split()
48 if len(N_k) != 2:
49 return False
50 N, k = map(int, N_k)
51 if not (1 <= N <= 100000):
52 return False
53 if not (1 <= k <= N):
54 return False
55 weights = list(map(int, lines[1:]))
56 if len(weights) != N:
57 return False
58 for w in weights:
59 if not (1 <= w <= 109):
60 return False
61 return True
62 except:
63 return False
64
65Note:
66- generate_test_input() must return a single adversarial input string, not a list.
67- Use cy.randint instead of cy.Integer().
68- The function should generate adversarial yet valid data fully inside, without parameters.
Listing 2: Adversarial Input Generator (CYaRon)
1Task:
2 Generate a challenging test input for the algorithm problem:
3 {problem_description}
4 Instructions:
5 - Focus on edge cases or scenarios that maximize the failure probability in faulty solutions.
6 - Due to the output length limit, you should generate a small-scale test input that is complete and valid.
7 - Output the test input directly, not code to generate it.
8 Output format:
9 ’’’plaintext
10 {test input}
11 ’’’
12 Think step by step.
@@ 8 Output format:
Appendix H Integrated Prompts for Algorithmic Problem Generation
1You are an expert competitive programmer.
2Ill provide you with two programming problems.
@@ 42. Design a new challenging problem that integrates them
5
6Output format(strictly follow):
7 ## Part 1: Original Problems and Solution Analysis
8 Step1: [Describe the steps of reasoning]
9 Step2: xxx
10 ...
11
12 ## Part 2: New Problem Description:
13 New_problem: [Describe the new problem clearly in natural language.]
14
15 Input Format: [Specify the input format]
16 Output Format: [Specify the output format]
17
18 ## Part 3: Example Test Cases
19 Input: [Input for test case 1]
20 Output: [Expected output for test case 1]
21 Input: [Input for test case 2]
22 Output: [Expected output for test case 2]
23
24 ## Part 4: Category
25 difficulty: [Easy/Medium/Hard]
26 tags: [tags of new problem, separated by commas.]
27 skills: [skills of new problem, separated by commas.]
28
29Note:
301. select the 1-3 most relevant skills from the given list, ranked by relevance.
@@ 22 Output: [Expected output for test case 2]
Listing 4: Prompts for Cross-type Fusion
1You are an expert competitive programmer.
2Ill provide you with one programming problem, its solution, and the key concepts they test.
3You need to:
@@ 41. Analyze its problem design approaches
52. Create a new variation question based on the original one
6
7Output format(strictly follow):
8 ## Part 1: Original Problems and Solution Analysis
9 Step1: [Describe the steps of reasoning]
10 Step2: xxx
11 ...
12
13 ## Part 2: New Problem Description:
14 New_problem: [Describe the new problem clearly in natural language.]
15
16 Input Format: [Specify the input format]
17 Output Format: [Specify the output format]
18
19 ## Part 3: Example Test Cases
20 Input: [Input for test case 1]
21 Output: [Expected output for test case 1]
22 Input: [Input for test case 2]
23 Output: [Expected output for test case 2]
24
25 ## Part 4: Category
26 difficulty: [Easy/Medium/Hard]
27 tags: [tags of new problem, separated by commas, referring to the tags of the original problems.]
28 skills: [skills of new problem, separated by commas.]
29
30Note:
311. select the 1-3 most relevant skills from the given list, ranked by relevance.
@@ 23 Output: [Expected output for test case 2]
Listing 5: Prompts for Single-problem Extension

### v2

@@ B.5.1 Setup and Notation
B.5.2 Systematic Bias (Identification and De-biasing)
@@
B.5.3 Random Error (Sampling Variability)
@@
By concavity of x(1x)x(1-x), (6) $\leq$ (8), so using (8) is conservative when the split is fixed.
@@
where $\underline{\mu},\overline{\mu}]$ is a $95\%$ CI for $\mu$. If only bounds are known ($\alpha\in[0,\alpha_{\max}]$, $q_{e}\in[q_{\min},q_{\max}]$), combine (5) with $\underline{\mu},\overline{\mu}]$ to obtain a conservative CI for $p$:
@@
B.5.4 Total Error Bound
@@
When $n$ is large, the $O(n^{-1/2})$ term vanishes and the total error is controlled by the bias ceiling $\alpha$. If $\alpha$ (and $q_{e}$) are known, report the de-biased estimate (4) with CI (10); this both removes the bias and shrinks the CI.
B.5.5 Stratified (Tag-wise) Contamination
@@ B.5.6 Numerical Illustration
B.5.7 Practical Safeguards
  • Audit and bound $\alpha$. Spot-check a random subsample to obtain an empirical upper bound $\alpha_{\max}$ with binomial CIs; report $p$ using (11).
@@ B.5.7 Practical Safeguards
  • De-bias when possible. If $(\alpha,q_{e})$ are fixed by design (e.g., known faulty items), publish the de-biased estimate (4) and its CI (10).
@@
  • Robust reporting. Alongside $\hat{\mu}$, report (i) de-biased $\tilde{p}$, (ii) contamination-aware CIs, and (iii) sensitivity bands under $(\alpha,q_{e})$ ranges as in (11).
@@
B.6 Complexity Bounds and Time Constraints
To illustrate the relationship between input scale, algorithmic complexity, and execution time, we analyze two approaches to a standard programming problem:
@@ B.6.1 Capacity Analysis
B.7 Integrated Prompts for Test Case Generation
Listing 1: Random Input Generator (CYaRon)
I will provide you with a programming problem description, and your task is to generate standardized test input samples using the CYaRon library.
You need to complete the following steps:
@@ B.7 Integrated Prompts for Test Case Generation
Output format (strictly follow)
Part 1: Parse Input Constraints
Specify the input constraints as described in the problem.
Part 2: Code for Test Input Generation
import cyaron as cy #cyaron version: 0.7.0
def generate_test_input():
# set parameters constraints that meet requirements (e.g. 1 <= N <= 300)
# Generate input using CYaRon
input_data = [
@@
Part 3: Code to Validate Test Input
def validate_test_input(input_string):
# Validation logic
return <boolean>
Note:
- cy.Integer() is not supported; it should be cy.randint.
- use cy.String.random instead of cy.String
- The function generate_test_input() should not accept any parameters. You need to generate the input entirely within the function.
- Generate code following the above format, without starting with ‘‘‘python or similar markers.
Listing 2: Adversarial Input Generator (CYaRon)
I will provide you with a programming problem description, and your task is to generate adversarial test input samples using the CYaRon library.
You need to complete the following steps:
@@
Output format (strictly follow):
Part 1: Parse Input Constraints
Specify the input constraints as described in the problem.
Part 2: Code for Test Input Generation
import cyaron as cy # cyaron version: 0.7.0
import random
def generate_test_input():
# set parameters constraints that meet requirements (e.g. 1 <= N <= 300)
# Randomly choose one adversarial strategy
strategy = random.choice(["equal_weights", "alternating_large_small", "large_ends"])
if strategy == "equal_weights":
# example: all weights are maximal
N = cy.randint(100000, 100000) # fix to worst-case size
max_weight = cy.randint(109, 109)
k = random.choice([1, N//2, N-1, N])
weights = [max_weight] * N
elif strategy == "alternating_large_small":
N = cy.randint(100000, 100000)
max_weight = cy.randint(109, 109)
k = random.choice([1, N//2, N-1, N])
weights = [max_weight if i%2 else 1 for i in range(N)]
else: # large_ends
N = cy.randint(100000, 100000)
max_weight = cy.randint(109, 109)
k = random.choice([1, N//2, N-1, N])
weights = [max_weight] + [1]*(N-2) + [max_weight]
# Build input string
input_lines = [f"{N} {k}"] + [str(w) for w in weights]
@@
Part 3: Code to Validate Test Input
def validate_test_input(input_string):
try:
@@
N_k = lines[0].split()
if len(N_k) != 2:
return False
N, k = map(int, N_k)
if not (1 <= N <= 100000):
return False
if not (1 <= k <= N):
return False
weights = list(map(int, lines[1:]))
if len(weights) != N:
return False
for w in weights:
if not (1 <= w <= 109):
return False
return True
except:
return False
Note:
- generate_test_input() must return a single adversarial input string, not a list.
- Use cy.randint instead of cy.Integer().
- The function should generate adversarial yet valid data fully inside, without parameters.
@@
Task:
Generate a challenging test input for the algorithm problem:
{problem_description}
Instructions:
- Focus on edge cases or scenarios that maximize the failure probability in faulty solutions.
- Due to the output length limit, you should generate a small-scale test input that is complete and valid.
- Output the test input directly, not code to generate it.
Output format:
’’’plaintext
{test input}
Think step by step.
B.8 Integrated Prompts for Algorithmic Problem Generation
Listing 4: Prompts for Compositional Variants
You are an expert competitive programmer.
Ill provide you with two programming problems.
If the problems test similar concepts (same-type fusion):
1. Analyze their problem design approaches
2. Create a new challenging problem testing the same concept(s).
If they test different concepts (cross-type fusion):
@@ 2. Design a new challenging problem that integrates them
you must choose one of the following variation strategies:
1. Sequential Combination: Concepts are loosely chained or simple spliced.
2. Deep Integration (Fusion): Concepts are merged in a non-trivial, deeply interconnected way.
Output format(strictly follow):
## Part 1: Original Problems and Solution Analysis
Step1: [Describe the steps of reasoning]
Step2: xxx
## Part 2: New Problem Description:
New_problem: [Describe the new problem clearly in natural language.]
Input Format: [Specify the input format]
Output Format: [Specify the output format]
## Part 3: Example Test Cases
Input: [Input for test case 1]
Output: [Expected output for test case 1]
Input: [Input for test case 2]
Output: [Expected output for test case 2]
## Part 4: Category
difficulty: [Easy/Medium/Hard]
tags: [tags of new problem, separated by commas.]
variation: [cross-type or same-type fusion; Sequential Combination or Deep Integration.]
Note:
1. Please generate a difficult and original question.
@@
Listing 5: Prompts for Atomic Variants
You are an expert competitive programmer.
Ill provide you with one programming problem, its solution, and the key concepts they test.
You need to:
@@ 1. Analyze its problem design approaches
2. Create a new variation question based on the orignal oneYou must choose **one** of the following variation strategies:
1. **Increase Data Scale:** Significantly increase the constraints, ensuring that an optimized algorithm can still solve the problem within a 5-second time limit.
2. **Rule Transformation (Preserving Core Algorithm):** Change the rules or constraints without affecting the fundamental algorithmic approach (e.g., in a "Climbing Stairs" problem, change steps from {1, 2} to {2, 3}).
Output format(strictly follow):
## Part 1: Original Problems and Solution Analysis
Step1: [Describe the steps of reasoning]
Step2: xxx
## Part 2: New Problem Description:
New_problem: [Describe the new problem clearly in natural language.]
Input Format: [Specify the input format]
Output Format: [Specify the output format]
## Part 3: Example Test Cases
Input: [Input for test case 1]
Output: [Expected output for test case 1]
Input: [Input for test case 2]
Output: [Expected output for test case 2]
## Part 4: Category
difficulty: [Easy/Medium/Hard]
tags: [tags of new problem, separated by commas, referring to the tags of the original problems.]
variation: [The type of variation used: Increase Data Scale or Rule Transformation]
Note:
@@
3. Select the variation type you are most confident in, aiming to use each of the two types with roughly equal probability over multiple interactions.
B.9 Integrated Prompts for Error Analysis
Listing 6: Prompts for Error Analysis
Task: You are an expert coding problem analyst. You will be given Original Problems, a New Variant Problem, Correct Code, and Wrong Code from different models. Follow the steps below.
#### Step 1: Problem Analysis
* Input: I will provide 1-2 original problems and 1 new variant problem.
* Action: Analyze the knowledge points and key testing points for both the original and new problems.
* Output Format:
> Original Problem 1: [Title]
> - Key Test Point: [Description]
> - Knowledge Point: e.g., Segment Tree, Dynamic Programming
> Original Problem 2: [Title] (if provided)
> - Key Test Point: [Description]
> - Knowledge Point: [Description]
> New Variant Problem: [Title]
> - Key Test Point: [Description]
> - Knowledge Point: [Description]
#### Step 2: Code Error Analysis
* Input: I will provide one correct code solution and several wrong code solutions from specific AI models.
* Action: For each wrong code, analyze its errors by selecting from the predefined Error Taxonomy below. You must identify all errors that the model made.
* Error Taxonomy (Choose from these 4 types):
1. Modeling Error: Selecting an incorrect or unnecessary algorithmic paradigm due to misinterpreting the problem variant (e.g., using greedy instead of DP).
2. Logic/Merge Bug: Wrong conditional logic or incorrect variable usage when merging subproblem results.
3. Indexing/Caching Bug: Incorrect cache sizing, array bounds, or off-by-one errors leading to crashes or incorrect outputs.
4. Complexity Error: Using an algorithm with unnecessarily high time/space complexity (e.g., O(n²) where O(n) suffices).
5. Others: Content is not code, or other implementation errors.
* Output Format for Each Wrong Model:
> Incorrect model: [model name]
> - Error Type: e.g, Modeling Error
> - Reason: [Concise explanation]
> - Erroneous Code Snippet: Code
> Incorrect model: [model name]
> - Error Type: e.g., Logic/Merge Bug, Complexity Error
> - Reason: [Concise explanation]
> - Erroneous Code Snippet: Code
Tips:
Error Types can only be selected from Modeling Error, Logic/Merge Bug, Indexing/Caching Bug, Complexity Error, Others.